1.Which of the following is equal to the sum of two consecutive even integers?
(A)144
(B)146
(C)147
(D)148
(E)149
2.If x^2+y^2=2xy,then x must equal
(A)-1
(B)0
(C)1
(D)-y
(E)y
Thank you so much for giving your time to answer my questions.

Answer by richard1234(7193)

1. If we let 2x and 2x+2 be two consecutive even integers, then their sum is 4x+2, which is 2 modulo 4. 146 is the only answer choice that is 2 mod 4, so B is correct.

2. We have x^2 - 2xy + y^2 = 0 --> (x - y)^2 = 0, x - y = 0. Hence x = y, answer choice E.